# Incidence estimates for tubes in complex space

### v2

@@ 6. Appendix
$\begin{split}\boldsymbol{v}_{1}&=\iota(\boldsymbol{u}_{1})=(1,0,0,0);\\ \boldsymbol{v}_{2}&=\iota(i\boldsymbol{u}_{1})=(0,1,0,0);\\ \boldsymbol{w}_{1}&=\iota(\boldsymbol{u}_{2})=\tfrac{1}{\sqrt{1+r^{2}}}(1,0,r\cos\zeta,r\sin\zeta);\\ \boldsymbol{w}_{2}&=\iota(i\boldsymbol{u}_{1})=\tfrac{1}{\sqrt{1+r^{2}}}(0,1,-r\sin\zeta,r\cos\zeta).\\ \end{split}$
@@
[…]
@@
[…]
This completes the proof of (18). For $n>2$, the orthogonal complements of $\iota(\ell_{1})$ and $\iota(\ell_{1})$ intersect in a subspace of dimension $42n-$ 4. Any slice of $\iota(N_{\delta}(\ell_{1})\cap\iota(N_{\delta}(\ell_{1}))$ by a translation of this subspace is a figure like that of the $C^{2}$ case. We can integrate the $4$-dimensional cross section volume over the common short directions to arrive at (19). ∎
@@ Acknowledgement
Throughout the course of this work, S.T. has been supported by an NSF GRFP fellowship and by NSF DMS-2037851. We would like to thank our advisor, Larry Guth, for introducing us to Kakeya and related problems and supporting us through the research. Also, we would like to thank Dominique Maldague for discussions about the initial formulation of the problem, and the anonymous referee(s) for their many concrete suggestions.

### v3

@@ 6. Appendix
[…]
@@
$\begin{split}|\iota(T_{1})\cap\iota(T_{2})|&=|N_{\delta}(V)\cap N_{\delta}(W)|=\int_{\mathbb{R}^{4}}1_{N_{\delta}(V)}1_{N_{\delta}(W)}\,dV\\ &=\int_{B_{\mathbb{R}^{2}}\left(\vec{0},\delta\right)}\left(\int_{\-\infty}^{\infty}\int_{-\infty}^{\infty}1_{N_{\delta}(W)}(\vec{x},\vec{y})\,dA_{\vec{x}}\right)\,dA_{\vec{y}}.\end{split}$ (20)
@@
$\begin{split}|\iota(T_{1})\cap\iota(T_{2})|&=\int_{B_{\mathbb{R}^{2}}\left(\vec{0},\delta\right)}\left(\int_{B_{\mathbb{R}^{2}}\left(\tfrac{1}{r}R_{\zeta}(\vec{y}),\tfrac{\delta\sqrt{1+r^{2}}}{r}\right)}1\,dA_{\vec{x}}\right)\,dA_{\vec{y}}\\ &\sim\int_{B_{\mathbb{R}^{2}}\left(\vec{0},\delta\right)}\frac{\delta^{2}(1+r^{2})}{r^{2}}\,dA_{\vec{y}}\,\sim\,\frac{\delta^{4}(1+r^{2})}{r^{2}}\,=\,\frac{\delta^{4}}{\sin^{2}\theta}.\end{split}$
This completes the proof of (18). For $n>2$, the orthogonal complements of $\iota(\ell_{1})$ and $\iota(\ell_{1})$ intersect in a subspace of dimension 2n42n-4. Any slice of $\iota(N_{\delta}(\ell_{1})\cap\iota(N_{\delta}(\ell_{1}))$ by a translation of this subspace is a figure like that of the $C^{2}$ case. We can integrate the $4$-dimensional cross section volume over the common short directions to arrive at (19). ∎
@@ Acknowledgement
Throughout the course of this work, S.T. has been supported by an NSF GRFP fellowship and by NSF DMS-2037851. We would like to thank our advisor, Larry Guth, for introducing us to Kakeya and related problems and supporting us through the research. Also, we would like to thank Dominique Maldague for her discussion about the initial formulation of the problem and her comments on an earlier draft, and the anonymous referee(s) for their many concrete suggestions.